# The Learning Approach to Games

### v2

@@ 2.1 Two player game example
In this section, we present a simple, repeatedly played two-player example. We demonstrate that even with a fixed one-step horizon, players can exhibit sophisticated dynamics. In this setting, both players learn transition costs $\hat{F}$ and the actions of their opponents $\hat{\Gamma}$, all within a fixed horizon $T=1$. Our primary argument is that formulating controls as an equilibrium does not adequately capture the dynamic strategies continually employed by the players. To address this shortcoming, we construct learning algorithms that capture these dynamic strategies. This example serves to illustrate why a more general framework is necessary for effectively modeling games, and it heuristically explores how the concept of equilibrium is inherently influenced by the learning process and the opponent itself.
@@
Let us point out that, with a one-step horizon, a unique Nash equilibrium exists that the first player is unwilling to play. Of course, by virtue of Folk’s theorem, any feasible outcome can be sustained in an infinitely repeated game. However, this result explicitly relies on the assumption that players are completely certain about their opponents’ future actions over an indefinite horizon. This strong assumption allows for almost any feasible value to be supported as an equilibrium, leaving the question of which outcome will be observed without a clear answer. Instead, we emphasize once again that the core element in games is the learning algorithms employed by the players. These algorithms naturally govern the (random) evolution of probability distributions over current and future actions, which in turn is sufficient to understand the evolution of the game, provided that players act rationally. Here, ”rational” does not necessarily mean ”clever” since players might easily adopt suboptimal learning algorithms. The design of robust learning algorithms is an area of significant and growing interest.
@@
Consider a fixed state and action space as $\mathbb{S}=\{0,1\}$ and $\mathbb{A}=[0,1]$. Suppose players are not learning the horizon and transition probabilities, that is, $\mathfrak{L}_{T}^{i},\mathfrak{L}_{p}^{i}$ are constants yielding $\hat{T}^{i}=1$ and $p^{i}(t,\vec{x},\vec{a},1)=a^{i}$. Initially, to state the rules of the game and to enable comparison with the Nash equilibrium, we assume players are not learning the state value or transition costs;
$\phi^{1}(t+1,X_{t+1}^{1},X_{t+1}^{2})=-{\bf 1}_{\left\{{X_{t+1}^{2}=1}\right\}},\hskip 30.00005ptF^{1}(t,X_{t}^{1},X_{t}^{2},a^{1},a^{2})=ca^{1},\ \ (c<1)$
$\phi^{2}(t+1,X_{t+1}^{1},X_{t+1}^{2})=-{\bf 1}_{\left\{{X_{t+1}^{1}\neq X_{t+1}^{2}}\right\}},\hskip 15.00002ptF^{2}=0$
@@
In words, the first player wants second player to move state 00, and the second player wants to be in the same state as the first player. However, the first player gains some by increasing the odds of moving to state 1. Now, costs of each player are
$J^{1}(t,\vec{x},\vec{a})=ca^{1}-\mathbb{P}^{t,\vec{x},\vec{a}}(X_{t+1}^{2}=1),\quad J^{2}(t,\vec{x},\vec{a})=-\mathbb{P}^{t,\vec{x},\vec{a}}(X_{t+1}^{1}\neq X_{t+1}^{2}),$ (2.15)
[…]
@@
From now on, as the one-step game does not depend on the current time and state, we will drop them from notations. Let us also mention that players makes decisions simultaneously. One could set it up as turn-based, but we aim to keep the game as symmetric as possible only excluding the cost structure.
@@
Note that there is a unique Nash equilibrium when the horizon is fixed at $1$, which is $\vec{a}=(1,1)$. Although exists, it is not necessarily useful for characterizing potential behaviors of the players. Once the first player fixes the action of the second player, the player becomes unaware of the intentions of the second player. Moreover, since players do not announce their strategies, searching for a Nash equilibrium with a larger horizon does not necessarily model this game either.
Now, we will construct a learning algorithm and numerically explore how the corresponding behaviors are evolving. To do so, we need to determine what players can observe. There are two immediate choices as at the time $t$: (i) players can observe the action taken by the other, or (ii) players can observe only the state. Notice that (i) is simpler because state can only be 00 or $1$, and does not carry all the information about the underlying action of the player. Thus, we consider (ii) from now on.
@@
Let us recap how the game is played in the perspective of the first player. First, we determine a probability $a^{1}$ for us to transition state $1$, and we get paid $ca^{1}$. Then, we loose \$$1$ if the other player ends up in state $1$. In the perspective of the second player, objective is to simply to follow the other player. We observe the past states of the other player and try to end up in the same state, and otherwise loose \$$1$. Due to the simplicity of the game, we can generate strategies that are expected to be observed. Starting from the second player, as she has a simpler cost structure;
@@
    Determine what is an acceptable level of noise in the observation of the other players state, which relies on the expectation of the player. If the recent observations are yielding a certain outcome (00 or $1$) consistent with the acceptable noise level, take the corresponding action. Else, start to explore other actions with rationale to penalize the noise.
In case of the first player, cost structure is slightly more intricate;
@@
    If the second player always appears at 00, explore larger actions to reduce cost (due to $-ca^{1}$). Keep increasing it until the second player starts to appear at state $1$ and overcomes the gains from $-ca^{1}$.
@@
    If the second player appears at $1$, which is costly, switch to actions that are not yet well explored. Keep searching until the second player appears at 00 regularly enough to keep realized cost consistent with expectations.
We remark that a straightforward $Q$-learning algorithm can be used to model players. However, $Q$-learning only models expected rewards for a given action, thus it converges to fixed actions and lacks the underlying details we aim to demonstrate. This underscores the same point: it is crucial to incorporate the design of the player to understand games.
@@ 2.1.1 Details of the Learning Algorithm
We now introduce the relevant parts of the framework specific for this problem. In general, each player models a transition cost $\hat{F}$ and an estimate $\hat{\Gamma}$ on the other player’s actions, relying on their observations held in the memory. Also, each player has an expectation defined as in (2.13).
Recall that observations of players are the realized states. That is, ${\cal E}^{1}={\cal E}^{2}=\{0,1\}$, and observations in (2.10) are depending on the realizations of the states. Here, $\Omega^{u}$ and $\mathbb{P}^{u}$ are determined by the random number generators that determines the transitions of states for the players at each round. We then set the observations ${\cal O}^{1},{\cal O}^{2}$ in equation (2.10) in an obvious manner. In the simulations, each player keeps a memory of a certain length, recording the realized states and costs.
Now, let $\big{\{}{\cal N}_{\Gamma}^{i,k}\big{\}}_{k=1}^{K}$ be the $i$’th player action networks where ${\cal N}_{\Gamma}^{i,k}:[0,1]\to[0,1]$. We then assing $\mathfrak{L}_{\Gamma}^{i}:\mathscr{E}^{i}\to([0,1]\mapsto{\cal P}([0,1]))$ in (2.11) as the empirical distribution formed by $\{{\cal N}_{\Gamma}^{i,k}\}_{k}$’s. That is,
@@
Similarly, we denote the cost networks as $\big{\{}{\cal N}_{F}^{i,\ell}\big{\}}_{\ell=1}^{K}$ [^8] We choose the same $K$ only for the notational simplicity., where ${\cal N}_{F}^{i,\ell}:[0,1]\to\mathbb{R}$. We then set $\hat{F}:\hat{\Omega}\times\mathbb{A}\to\mathbb{R}$ in (2.5) as
$\hat{F}^{1}(\ell,\hat{\omega}^{\prime},a^{1})=-ca^{1}+{\cal N}_{F}^{1,\ell}(\hat{\omega}^{\prime})(a^{1}),\quad\hat{F}^{2}(\ell,\hat{\omega}^{\prime},a^{2})={\cal N}_{F}^{2,\ell}(\hat{\omega}^{\prime})(a^{2})$
We identify $(\ell,\hat{\omega}^{\prime})\in\hat{\Omega}=\{1,\dots,K\}\times\hat{\Omega}^{\prime}$ where $\hat{\mathbb{P}}$ assigns the first marginal as uniform distribution over $\{1,\dots,K\}$. [^9] As in the case of action networks, this is only for simplicity. One might assign and evolve weights corresponding to networks, and capture more vibrant dynamics if the game is more sophisticated. For example, one might keep a subset of networks as trusted ones (high weights), and let other networks explore more wildly (low weights). Each network ${\cal N}_{F}^{i,\ell}$ is further random by the virtue of dropout layers. Keeping the networks always in the training mode, one generates a random function with positive dropout probabilities, and $\hat{\Omega}^{\prime}$ characterizes this. Here,
@@
There are two objectives cost networks are training for: (i) there is an expected cost coming from the predictions of action networks, which is (2.15) integrated as in (2.6). If action networks are not perfect, expected cost will not match the observed expected cost. Cost networks are training to close this gap, by relying on a cost memory similar to action memory. (ii): players have expectations over what is best possible as introduced in (2.13), which we took as constant for simplicity. Cost networks are also trained such that the players do not get desperate. In the simulations, if the first player ends up with networks ${\cal N}_{\Gamma}^{1,k}$’s taking values close to $1$, independent of the action, they start to play the Nash equilibrium $(1,1)$. That is, player $1$ gets desperate, and then adjusts the random component of cost to start exploring other actions.
@@
$J^{1}(\ell,\hat{\omega}^{\prime},a^{1})=\frac{1}{K}\sum_{k=1}^{K}{\cal N}_{\Gamma}^{1,k}(a^{1})-ca^{1}+{\cal N}_{J}^{1,\ell}(\hat{\omega}^{\prime})(a^{1}),\ \text{ and}$
$J^{2}(\ell,\hat{\omega}^{\prime},a^{2})=\frac{1}{K}\sum_{k=1}^{K}{\cal N}_{\Gamma}^{1,k}(a^{2})+a^{2}(1-2{\cal N}_{\Gamma}^{1,k}(a^{2}))+{\cal N}_{J}^{1,\ell}(\hat{\omega}^{\prime})(a^{2})$
To draw an action from (2.8), players randomly choose one cost network $\ell\in\{1,\dots,K\}$, and observe one realization $\hat{\omega}^{\prime}$ coming from dropout layers. Then, they simply minimize $J^{i}$ over $a^{i}$, that is $\hat{\pi}$ yields deterministic $\varepsilon$-optimal action, and play that action.
Before discussing the simulation results, let us mention that each parameter of networks of course plays a significant role, and we coarsely tuned them by hand to obtain simulations matching with expectations. Many of such parameters are taken as constant. Changing to different constants might of course yield poor results. On the other hand, generalizing them will improve the sophistication of agents strategies. Besides the network parameters, there are more structural parameters too. For example, what players are expecting as the best possible cost also changes the characteristics of actions. Especially if the first player is expecting much better than what is realistically possible, exploration gets out of hand and they don’t converge anywhere. For an another example, if the memory is very long and not forgetting, then the first player starts to get advantage over the second player, as the second player becomes fixed after a while and ”thinks” that the other player will still frequently move to state 00. The point we are aiming to convey here is that the learning algorithms of players are crucial to characterize what is going to be realized in reality, and the evolution of ${}^{n}\gamma_{\mathfrak{L}}^{i}\in{\cal P}([0,1])$ might be dynamic even in the simplest settings.
@@
Figure 1: Actions of Player $1$ (red) and Player $2$ (blue) over $1000$ games, demonstrating the impact of varying the incentive parameter $c$ and expectation $B^{1}$ for Player $1$, while Player $2$’s parameters remain constant. Each subplot shows how different incentives and expectations influence Player $1$’s strategy and interaction dynamics in this toy problem.
@@
In subplot (a), Player 1 has a somewhat large incentive ($c=3/10$) to take larger actions, aiming for a reward of $B^{1}=1/10$. Thus, playing close to $(0,0)$ does not suffice, and Player 1 searches for higher rewards, leading to frequent changes between different phases. Notice that as soon as $(a^{1},a^{2})$ approaches $(1,1)$, Player 1 begins to explore and pushes the game back towards $(0,0)$. In subplot (b), the incentive is much smaller ($c=1/20$) and the expectation of Player 1 is decreased to $B^{1}=-1/10$. Consequently, Player 1 stays close to $(0,0)$ for longer, before starting to think that Player 2 will always choose action 00. In subplots (c) and (d), the incentive for Player 1 is really high, making deviations from $(1,1)$ unnecessary. Specifically, in subplot (d), Player 1 is satisfied with a reward of 00, maintaining $(1,1)$ almost always. Conversely, in subplot (c) where Player 1 expects to get an unrealistic reward of $1$, exploration by Player 1 leads to worse results for both.
We refer to the supplementary online repository [20] for the animation showing the cost networks, action networks, and other observables in each case.

### v4

@@ 3.1 Two player game example
In this section, we present a simple, repeatedly played two-player example. We demonstrate that even with a fixed one-step horizon, players can exhibit non-stationary dynamics. In this setting, both players learn transition costs $\hat{F}$ and the actions of their opponents $\hat{\Gamma}$, all within a fixed horizon $T=1$. Our central argument is that formulating controls as an equilibrium does not adequately capture the dynamic strategies continually employed by the players. To address this shortcoming, we construct learning algorithms that capture these dynamic strategies. This example illustrates why a more general framework is necessary for effectively modeling games, and it also heuristically highlights how the concept of equilibrium is inherently shaped by the learning process and by the opponents themselves.
Consider fixed state and action spaces given by $\mathbb{S}=\{0,1\}$ and $\mathbb{A}=[0,1]$. Players’ actions determine their transition probabilities at each step, and they can only observe each other’s state. The first player loses $\mathdollar 1$ if the second player appears in state $1$ but gains by increasing their own transition probability to state $1$. The second player loses $\mathdollar 1$ if they are not in the same state.  [^7] We mention their gains and losses in dollar amounts, to relate easily to a potential game we might play in real life.
@@
Let us note that, with a one-step horizon, there exists a unique Nash equilibrium that the first player is unwilling to play. Of course, by virtue of Folk’s theorem, any feasible outcome can be sustained in an infinitely repeated game. However, this result explicitly relies on the assumption that players are certain about their opponents’ future actions over an indefinite horizon. This strong assumption allows for almost any feasible value to be supported as an equilibrium, leaving the question of which outcome will be observed without a clear answer. Moreover, since players do not announce their strategies, searching for a Nash equilibrium with a larger horizon does not necessarily model this game either. Such a search represents our external attempts to formalize the players’ incentives. Instead, we emphasize once again that the core element in games is the learning algorithms employed by the players. These algorithms naturally govern the (random) evolution of probability distributions over actions, which in turn is sufficient to understand the evolution of the game.
@@
Now, let us formally state the game. Consider fixed state and action spaces given by $\mathbb{S}=\{0,1\}$ and $\mathbb{A}=[0,1]$. Suppose the players are not learning the horizon or transition probabilities, that is, $\mathfrak{L}_{T}^{i},\mathfrak{L}_{p}^{i}$ are constants yielding $\hat{T}^{i}=1$ and $p^{i}(t,\vec{x},\vec{a},1)=a^{i}$. Initially, to specify the rules of the game and enable comparison with the Nash equilibrium, we assume that the players are not learning the state value or transition costs;
$\phi^{1}(t+1,X_{t+1}^{1},X_{t+1}^{2})=-{\bf 1}_{\left\{{X_{t+1}^{2}=1}\right\}},\hskip 20.00003ptF^{1}(t,X_{t}^{1},X_{t}^{2},a^{1},a^{2})=ca^{1},$
$\phi^{2}(t+1,X_{t+1}^{1},X_{t+1}^{2})=-{\bf 1}_{\left\{{X_{t+1}^{1}\neq X_{t+1}^{2}}\right\}},\hskip 5.0ptF^{2}=0$
@@
As mentioned, the first player wants the second player to move state 0, and the second player wants to be in the same state as the first. However, since $c>0$, the first player gains by increasing the oods of moving to state $1$. Now, costs of each player are
$J^{1}(t,\vec{x},\vec{a})=ca^{1}-\mathbb{P}^{t,\vec{x},\vec{a}}(X_{t+1}^{2}=1),\ \ J^{2}(t,\vec{x},\vec{a})=-\mathbb{P}^{t,\vec{x},\vec{a}}(X_{t+1}^{1}\neq X_{t+1}^{2}),$ (3.16)
$\mathbb{P}^{t,\vec{x},\vec{a}}(X_{t+1}^{2}=1)=a^{2},\hskip 20.00003pt\mathbb{P}^{t,\vec{x},\vec{a}}(X_{t+1}^{1}\neq X_{t+1}^{2})=a^{1}+a^{2}(1-2a^{1})$
@@
From now on, since the one-step game does not depend on the current time or state, we will drop them from notations. Let us also note that players make their decisions simultaneously. One could instead formulate the game as turn-based, but we aim to keep it as symmetric as possible, while excluding only the cost structure.
@@
Note that when the horizon is fixed at $1$, there exists a unique Nash equilibrium given by $\vec{a}=(1,1)$. Although such an equilibrium exists, it is not necessarily useful for characterizing the potential behaviors of the players. Once the first player fixes the action of the second player, they lose awareness of the latter’s underlying intentions.
@@
Let us recap how the game is played from the perspective of the first player. First, we determine a probability $a^{1}$ of transitioning to state $1$, and receive a payoff $ca^{1}$. Then, we lose \$11 if the other player ends up in state $1$. From the perspective of the second player, objective is simply to follow the other player. We observe the past states of the other and attempt to end up in the same state, losing \$11 otherwise. Due to the simplicity of the game, behaviors that are expected to recur over time can be generated. Starting with the second player, whose cost structure is simpler:
@@
    Determine an acceptable level of noise in observing the other player’s state, based on expectations. If recent observations consistently yield a particular outcome (0 or $1$) within the acceptable noise level, take the corresponding action. Otherwise, begin exploring other actions, with rationale to penalize the noise.
For the first player, the cost structure is slightly more intricate;
@@
    If the second player consistently appears in state 0, explore larger actions to reduce the cost (due to $ca^{1}$). Continue increasing it until the second player begins to appear in state $1$ frequently enough to offset these gains.
@@
    If the second player appears at $1$, which is costly, switch to actions that are not used recently. Continue exploring until the second player reappears in state 0 regularly enough to keep the realized cost consistent with expectations.
In the next section, we construct a learning algorithm and numerically explore how the corresponding behaviors evolve. We remark that a straightforward $Q$-learning algorithm can be used to model the players. However, $Q$-learning models only expected rewards for each action, it therefore converges to fixed actions and lacks the underlying dynamics we aim to demonstrate. This underscores the same point: it is crucial to incorporate the design of the player to understand games.
@@ 3.1.1 Details of the Learning Algorithm
We now introduce the relevant parts of the framework specific for this problem. In general, each player models a transition cost $\hat{F}$ and an estimate $\hat{\Gamma}$ on the other player’s actions, relying on their observations held in the memory. Also, each player has an expectation $\hat{B}$ as in (3.14), taken as a constant for simplicity.
Recall that observations of players are the realized states. That is, ${\cal E}^{1}={\cal E}^{2}=\mathbb{S}$, and observations in (3.10) are depending on the realizations of the states. Here, $\Omega^{u}$ and $\mathbb{P}^{u}$ are determined by the random number generators that determines the transitions of states for the players at each round. We then set the observations ${\cal O}^{1},{\cal O}^{2}$ in equation (3.10) in an obvious manner. In the simulations, each player keeps a memory of a certain length, recording the realized states and costs.
Now, let $\big\{{\cal N}_{\Gamma}^{i,k}\big\}_{k=1}^{K}$ be the $i$’th player simple feed-forward networks where ${\cal N}_{\Gamma}^{i,k}:\mathbb{A}\to\mathbb{A}$. We then assing $\mathfrak{L}_{\Gamma}^{i}:\mathscr{E}^{i}\to(\mathbb{A}\mapsto{\cal P}(\mathbb{A}))$ in (LABEL:eqn:learningalgorithms) as the empirical distribution formed by $\{{\cal N}_{\Gamma}^{i,k}\}_{k}$’s. That is,
@@
Similarly, we denote $\big\{{\cal N}_{F}^{i,\ell}\big\}_{\ell=1}^{K}$, where ${\cal N}_{F}^{i,\ell}:[0,1]\to\mathbb{R}$. We then set $\hat{F}:\hat{\Omega}\times\mathbb{A}\to\mathbb{R}$ in (3.5) as
$\hat{F}^{1}(\ell,\hat{\omega}^{\prime},a^{1})=ca^{1}+{\cal N}_{F}^{1,\ell}(\hat{\omega}^{\prime})(a^{1}),\quad\hat{F}^{2}(\ell,\hat{\omega}^{\prime},a^{2})={\cal N}_{F}^{2,\ell}(\hat{\omega}^{\prime})(a^{2})$
We identify $(\ell,\hat{\omega}^{\prime})\in\hat{\Omega}=\{1,\dots,K\}\times\hat{\Omega}^{\prime}$ where $\hat{\mathbb{P}}$ assigns the first marginal as uniform distribution over $\{1,\dots,K\}$. [^8] As in the case of action networks, this is only for simplicity. One might assign and evolve weights corresponding to networks, and capture more vibrant dynamics if the game is more sophisticated. For example, one might keep a subset of networks as trusted ones (high weights), and let other networks explore more wildly (low weights). Each network ${\cal N}_{F}^{i,\ell}$ is further random by the virtue of dropout layers. Keeping the networks always in the training mode, one generates a random function with positive dropout probabilities, and $\hat{\Omega}^{\prime}$ abstracts this. Here,
@@
There are two objectives cost networks are training for: (i) there is an expected cost coming from the predictions of action networks, which is (3.16) integrated as in (3.6). If action networks are not perfect, expected cost will not match the observed expected cost. Cost networks are training to close this gap, by relying on costs in the memory. (ii): players have expectations over what is best possible as introduced in (3.14), which we took as constant for simplicity. Cost networks are also trained such that the players do not get desperate. For example, if the first player ends up with networks ${\cal N}_{\Gamma}^{1,k}$’s taking values close to $1$, independent of the action, they start to play the Nash equilibrium $(1,1)$. That is, player $1$ gets desperate, and then adjusts the random component of cost to increase values of other actions towards $\hat{B}^{1}$ to start exploring them.
@@
$J^{1}(\ell,\hat{\omega}^{\prime},a^{1})=\frac{1}{K}\sum_{k=1}^{K}{\cal N}_{\Gamma}^{1,k}(a^{1})+ca^{1}+{\cal N}_{F}^{1,\ell}(\hat{\omega}^{\prime})(a^{1}),\ \text{ and}$
$J^{2}(\ell,\hat{\omega}^{\prime},a^{2})=\frac{1}{K}\sum_{k=1}^{K}{\cal N}_{\Gamma}^{2,k}(a^{2})+a^{2}(1-2{\cal N}_{\Gamma}^{2,k}(a^{2}))+{\cal N}_{F}^{2,\ell}(\hat{\omega}^{\prime})(a^{2})$
To draw an action from (3.9), players uniformly choose one cost network $\ell\in\{1,\dots,K\}$, and observe one realization $\hat{\omega}^{\prime}$ coming from dropout layers. Then, ${\Upsilon}^{i}=\hat{\pi}^{i}$ simply minimizes $J^{i}$ over $a^{i}$ yielding $\varepsilon$-optimal action, and plays it.
Before discussing the simulation results, let us mention that each parameter of networks of course plays a significant role, and we coarsely tuned them by hand to obtain simulations matching with expectations. Many of such parameters are taken as constant. Changing to different constants might of course yield poor results. On the other hand, generalizing them will improve the sophistication of agents strategies. Besides the network parameters, there are more structural parameters too. For example, what players are expecting as the best possible cost also changes the characteristics of actions. Especially if the first player is expecting much better than what is realistically possible, exploration gets out of hand. For an another example, if the memory is very long and not forgetting, then the first player starts to get advantage over the second player, as the second player becomes fixed after a while and estimates that the other player will still frequently move to state 0. The point we are aiming to convey here is that the learning algorithms of players are crucial to characterize what is going to be realized in reality.
@@
Figure 1: Actions of Player $1$ (red) and Player $2$ (blue) over 10001000 games, demonstrating the impact of varying the incentive parameter $c$ and expectation $B^{1}$ for Player $1$, while Player $2$’s parameters remain constant. Each subplot shows how different incentives and expectations influence Player $1$’s strategy and interaction dynamics in this toy problem.
@@
In subplot (a), Player 1 has a somewhat large incentive ($c=3/10$) to take larger actions, aiming for a reward of $B^{1}=1/10$. Thus, playing close to $(0,0)$ does not suffice, and Player 1 searches for higher rewards, leading to frequent changes between different phases. Notice that as soon as $(a^{1},a^{2})$ approaches $(1,1)$, Player 1 begins to explore and pushes the game back towards $(0,0)$. In subplot (b), the incentive is much smaller ($c=1/20$) and the expectation of Player 1 is decreased to $B^{1}=-1/10$. Consequently, Player 1 stays close to $(0,0)$ for longer, before starting to think that Player 2 will always choose action 0. In subplots (c) and (d), the incentive for Player 1 is really high $(c=1)$, making deviations from $(1,1)$ unnecessary. Specifically, in subplot (d), Player 1 is satisfied with a reward of 0, maintaining $(1,1)$ almost always. Conversely, in subplot (c) where Player 1 expects to get an unrealistic reward of $1$, which requires Player $1$ to play $1$ while Player $2$ plays 0, exploration by Player 1 leads to worse results for both.
We refer to the supplementary online repository [42] for the animation showing the cost networks, action networks, and other observables in each case. We point out that these estimations are not converging networks, instead dynamic and yielding repeating patterns of behaviors.